OpenStudy (anonymous):

a plane flies in a straight line path after taking off, and it ends up 210 km farther east and 90.0 km farther north, relative to where it started. What is the plane's displacement?

OpenStudy (radar):



OpenStudy (anonymous):

or the first one use Pythagorean. distance^2= 210^2 + 90^2 =228.47km, about 230km For the second use: Inverse-tangent of 90/200 =24.23 degrees, about 24 degrees

OpenStudy (radar):

Typo 90/200 should read 90/210